A lot consists of 10 good articles, 4 articles with minor defects and 2 with major defects. Two articles are chosen at random from the lot (without replacement) Find the probability that: (a) both are good(b) both has major defects, (c) at least 1 is good (d) neither has major defects, (e) neither is good


1
Expert's answer
2021-01-27T02:12:34-0500

There are total 16 articles out of which 10 good articles, 4 articles with minor defects and 2 articles with major defects.

Two articles can be chosen in "{^{16}C_2}" ways. Therefore the sample space contains "^{16}C_2" elements.

(a) Let "A" be an event such that both the articles are good.

Then the number of elements in favour of "A" is "{^{10}C_2}".

"\therefore" "P(A)=\frac{^{10}C_2}{^{16}C_2}=\frac{45}{120}=\frac{3}{8}"

(b) Let "B" be an event such that both the articles have major defects

Then the number of elements in favour of "B" is "{^{2}C_2}"

"\therefore" "P(B)=\frac{^{2}C_2}{^{16}C_2}=\frac{1}{120}"


(c) Let "C" be an event such that at least 1 article is good. Then two cases arise.

Case 1: It will be 1 good article and 1 article with major or minor defects

Case 2: 2 articles are good.

For the first case, 1 good article and 1 article with major or minor defects can be chosen in "{^{10}C_1}\u00d7{^{6}C_1}" ways .

For second case , 2 good articles can be chosen in "{^{10}C_2}" ways.

So total number of elements in favour of "C" is "({^{10}C_1}\u00d7{^{6}C_1})+{^{10}C_2}=(60+45)=105"

"\therefore" "P(C)=\frac {105}{120}=\frac{7}{8}"


(d) Let "D" be an event such that neither has major defects. Then articles will be either good or have minor defects.

Then the number of elements in favour of "D" is "{^{14}C_2}"

"\therefore" "P(D)=\frac {^{14}C_2}{^{16}C_2}=\frac{91}{120}"


(e) Let "E" be an event such that neither is a good article. Then articles will be either major or minor defects.

Then the number of elements in favour of "E" is "{^6C_2}"

"\therefore" "P(E)=\frac{^6C_2}{^{16}C_2}=\frac{15}{120}=\frac{1}{8}"